1.-Determine the payment to amortize the debt. (Round your answer to the nearest cent.) Quarterly payments on $15,500 at 3.7% for 6 years

$

2.- A MasterCard statement shows a balance of $510 at 13.5% compounded monthly. What monthly payment will pay off this debt in 1 year 10 months? (Round your answer to the nearest cent.)

$

3.- Find the unpaid balance on the debt. (Round your answer to the nearest cent.) After 7 years of monthly payments on $180,000 at 3% for 25 years.

$
Answer #1

1) PV = 15,500, yield = 3.7% paying quarterly, t = 6 years find yield to amortize the payment

15,500 = x/(1+0.037/4)+..........x/(1+0.037/4)^24

x = 723.14

2)

513 = x/(1+0.135/12)+........x/(1+0.135/12)^13

x = 42.63

3)

180,000 = x/(1+0.03/12)+........x/(1+0.03/12)^300

x = 853.58

after 7 years unpaid amount = 853.58/(1+0.03/12)+.......853.58/(1+0.03/12)^216

= 142,328.97

converting in present value = 142,328.97/(1+0.03)^7

= 115,726.47